OpenStudy (anonymous):

Solve for R 1/R= 1/R1 + 1/R2

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE 1}{\LARGE R} =\frac{\LARGE 1}{\LARGE R_1}+\frac{\LARGE 1}{\LARGE R_2} }\) like this, correct?

OpenStudy (zarkon):

reciprocate both sides

OpenStudy (anonymous):

@SolomonZelman yes

OpenStudy (anonymous):

@Zarkon can you show me how you would solve it? I did it and it got messy

OpenStudy (anonymous):

@Zarkong after I reciprocate I get R = R/R1 + R/R2

OpenStudy (zarkon):

\[R=\frac{1}{\frac{1}{R_1}+\frac{1}{R_2}}=\frac{R_1R_2}{R_1+R_2}\]